OpenStudy (anonymous):

find functions of f and g that show that is incorrect: d f(x) (d/dx)f(x) -- --- = --------- dx g(x) (d/dx)g(x)

OpenStudy (anonymous):

f[x]=sin[x] g[x]=cos[x]

OpenStudy (anonymous):

d/dx sin x = cos x.......................[1] d/dx cos x =- sin x ....................[2] [1]/[2] = -cot[x] \[d/dx (tan x) = sec^2(x)\] \[sec^2(x) \neq -cot(x) \]

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

yeah, that makes sense. cheers, cyter. Which other pair could disprove this?

OpenStudy (anonymous):

f[x]=2x g[x]=3x f'[x]/g'[x] = 2/3 d/dx(f[x]/gx)= d/dx(2/3) = 0 \[0\neq \frac{2}{3}\]